This Question: 5 pts 9 of 40 (8 complete) Factor v3 +0.008 Select the correct choice below and, if necessary, fill in the answer box within your choice. A. +0008- (Factor completely. Use integers or decimals for any numbers in the expression.) B. The polynomial is prime. nter vour answerís)
Answer #1

Given polynomial is:

+ 0.008

We can write it as:

y^3+0.008= (y)^3+(0.2)^3

As we know a^3+b^3 = (a+b)(a^2-ab+b^2)

So y^3+(0.2)^3 = (y+0.2)(y^2-y*0.2+(0.2)^2)

y^3+(0.2)^3 = (y+0.2)(y^2-0.2y+0.04)